# Concentration Inequalities for Branching Random Walks

### v4

@@ Abstract
While classical concentration inequalities are typically restricted to two special cases—independence and martingale difference sequences—we extend concentration inequalities to a much broader class of stochastic processes by relaxing these foundational conditions. Specifically, heuristically and in the language of calculus, while independence and the martingale difference property correspond to
@@
respectively, we relax these conditions to
@@
Furthermore, concentration inequalities are established for branching random walks.
@@ 1.1 Concentration inequalities
Let $(S_{i})_{i=1,2,...}$ be a real-valued random process, one-dimensional random walk. $S_{n}$ can be formulated as $S_{n}=X_{1}+X_{2}+...+X_{n}$; $X_{i}=S_{i}-S_{i-1}$. If some conditions are satisfied, among which independence is crucial, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT). For $S_{n}$ beyond the neighborhood, i.e., $|S_{n}-E(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [23]), also referred as Chernoff bound,
$Pr(|S_{n}-E(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
@@
However, for a long time, there had been no significant improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [40]. The reason might be two folds. First, there lacked motivation. Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results concealed the essential reason why the inequalities hold. In early 2000s [37] relaxed the boundedness to finite exponential moments; i.e. for a $\delta>0,E(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-E(S_{n})|\geq\lambda|)\leq e^{-\sim\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$ — not Chernoff bound. Later, motivated by the study of the free energy of a directed polymer in a random environment, [32] gave Chernoff bound under the hypothesis that, for a $\delta>0,E(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$). Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [25]. No attempt is known to get rid of martingale condition. The first major contribution of this paper is in this direction.
1.2 Notation and contribution
@@
We let $\mathbb{E}_{i}[\,\cdot\,]$ to abbreviate $\mathbb{E}[\,\cdot\mid u_{i}\,]$. Note that $\mathbb{E}_{i}[X_{t}]$ is a function of both $t$ and past state $(u_{i})_{i\leq t}$. To connect the calculus notation with the underlying stochastic process, we define
@@ 1.4 Motivation and contribution
In algorithm analysis and random graphs study, which are described as graph evolution stochastic process (see e.g. [34], [42], [43], [44], [45] and references therein), the differential equation method (DEM) started in 80’s. Recently DEM has obtained some important results and received increasingly attention. The procedure of DEM consists of (a) verifying that the random parameters concerned concentrate around their means, and (b) finding solutions to the differential equations satisfied the means. Concentration inequalities (e.g. Heoffding-Azuma inequalities) are used to prove concentration phenomena in DEM.
@@
2 An extension of Heoffding-Azuma inequality
For warming up, let’s present a “downgraded” proof for Azuma’s inequality. Possibly this proof could be found in the Azuma’s trash can. Let’s first introduce a simple and elegant inequality due to [32] and [41]. Given here is the version from [41] which improves [32].
@@ 2 An extension of Heoffding-Azuma inequality
$Pr(|S_{n}-ES_{n}|\geq\lambda)\leq 2\exp\Big(-\frac{\lambda^{2}}{2nc^{2}}\Big)$
@@
which is exponentially small under the choice $t\sim n^{-1/2}$, even for $\lambda\ll n$.
Proof (Downgraded version of Azuma’s inequality).
We only consider $\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$, as the proof of the reverse $\mathbb{P}(S_{n}-\mathbb{E}S_{n}\leq-\lambda)$ is exactly the same. By Markov’s inequality, inductive application of Lemma 2.1 and optimization of $t$, we have
$\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$ $\leq e^{-t\lambda}\cdot E\{e^{tS_{n-1}}\cdot\mathbb{E}(e^{tX_{n}}\mid X_{1},...,X_{n-1})\}$
@@
$\leq e^{-t\lambda}\cdot\mathbb{E}\{e^{tS_{n-1}}\cdot e^{t^{2}e^{c}}\}~~~~~~~\mbox{(}|X_{i}|\leq c\mbox{)}$
@@
The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [32] and [41]) without compromising strength of results; from the practical point of view, the difference between $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ and $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$ is immaterial – both are square-exponential decay. In addition, it is demonstrated that nifty techniques are not necessary for obtaining concentration inequalities.
Next, we will show what else can make the concentration inequality hold, other than independence and martingale difference. To this end, define Doob’s (or McDiarmid’s) martingale
@@
$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\quad$
$E_{i-1}(d_{i})=0$. Notice
@@
Note, $\mathbb{E}_{i}(X_{\ell})=\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
@@
With this notation,
@@
$\leq\mathbb{E}_{X_{i}^{\prime}}\Big(\,\bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\bigl|\,\Big)$
@@
and $\mathbb{E}[\,e^{\delta A|X_{i}|}\mid u_{i-1}\,]\leq K$ (exponential moment existence) for a constant $\delta>0$ , $i=1,2,$..., then for $n\leq M$
$\Pr\big(\,|\,S_{n}-\mathbb{E}S_{n}\,|\geq\lambda\,\big)\leq 2\exp{\Big(\!-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\Big)}$ (2.9)
Obviously, Poisson distribution $\text{Poisson}(\lambda_{i})$ meets the exponential moment condition.
@@
$\mathbb{E}_{i-1}\exp\left(\delta|d_{i}|\right)=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
@@
$\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)\leq\exp\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right).$
For the other half side of the concentration inequality, we refer the fact that Lemma 2.1 holds for negative $t$ as well. The proof is exactly the same as above, which we omit. ∎
@@
    where
  • The condition of Lipschitz continuity (2.10) may be rephrased “uniqueness of solution”; i.e. small difference in initial position can not yield big difference in the trajectory.
@@
  • The condition of Lipschitz continuity (2.10) may be rephrased “uniqueness of solution”; i.e. small difference in initial position can not yield big difference in the trajectory.
@@
Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on. We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the region in question is distant enough from singularities).
@@
If $E\big(e^{\delta A|X_{i}|}\mid u_{i-1}\big)\leq K$ (existence of an exponential moment), then in the region where $\mu_{i}$ and $\nu_{i}$ are smooth, the concentration inequality holds.
@@
$\frac{\partial\mu(t,u)}{\partial u},\quad\frac{\partial^{2}\mu(t,u)}{\partial u^{2}},\quad\frac{\partial\nu(t,u)}{\partial u},\cdots$
are uniformly bounded in the region of smoothness (i.e., away from singularities), in addition to the exponential moment condition. These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE). From the PDE we have every partial derivatives of $\bar{u}$, and $u$ is close to $\bar{u}$ with exponentially rare and negligible exceptions. Note, the condition (2.10) says that small changes in distant past generations won’t lead to violently change in the current step size. Theorem 2.2 tells us, in terms of PDE of $\bar{u}$, what local behavior of $u$ can meet (2.10), for the concentration inequality to hold.
Proof.
We just need to prove Lipschitz continuity (2.10). The proof is proceeded by induction. The base case holds, i.e. $|E(X_{i+1}|u_{i})-E(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
@@
the following recurrence equation about $\mathrm{Var}_{n+1}$ holds.
@@
Now let’s estimate $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$. From (2.13) and (2.12), we obtain
@@
Given above, it is easy to check that, in the region where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$. Further, by induction, $\frac{d\mathrm{Var}_{n+1}}{du}$ exists and its order is $O(1/N)$.
@@ 3 Concentration inequalities in the context of BRW
$\mathbb{E}^{\!{}^{(M)}}[\,\cdot\,]=\frac{\displaystyle\int(\,\cdot\,)\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}{\displaystyle\int\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}$
@@
$Pr[\ |S_{n}-\mathbb{E}^{{(M)}}(S_{n})\mid\geq\lambda]\leq 2\exp{\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)},\quad\text{$n\leq M$ }$ (3.7)
@@
If $E(e^{\delta A|X_{i}|}\mid u_{i-1})\leq K$, then in the area where $m_{i}(u)$, $\mu_{i}(u)$ and $\nu_{i}(u)$ are smooth, the concentration inequality holds.
@@
$\overline{X}_{n+1}^{(k)}=\mathbb{E}_{n}^{(k)}(X_{n+1}),\quad f_{n+1}^{(k-1)}(u_{n+1})=\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right];$
@@
which is smooth in $u_{n}$. Observe that the term $\frac{\overline{X}_{n+1}^{(k)}}{N}{f_{n+1}^{(k-1)}}^{\prime}(u_{n})$ involves no second order derivatives of $f_{n+2}^{(k-2)}$ which is absorbed into the $O(N^{-2})$ term. Thus, we have proved that for any $k$, $\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]$ is smooth in $u_{n}$.
@@
In the situation of bounded $X_{i}$, a particle always stays around the mean trajectory as long as its birthplace $u_{i-1}$ is near that trajectory. In the case of unboundedness, however, a particle may move into regions near singularities (e.g., critical points in K–SAT) where $u_{i}$ is not smooth. In such regions, $\mathbb{E}(e^{\delta Xi})<K$ fails, among others. The good news is, because concentration inequality holds in the neighborhood area of the mean trajectory—is at least $\epsilon$ far away from the singularity—the outliers account for a negligible proportion of the population and can be ignored. Crucially, pruning these rare extremal paths does not weaken the concentration. The following theorem formalizes this observation.
@@
Let $\epsilon$ be small compared with $\delta$, $N^{1/2}\ll\lambda<\epsilon N$, and as before,
@@
$G_{i-1}:=\Big(\bar{u}_{i-1}-\frac{\lambda}{N},\bar{u}_{i-1}+\frac{\lambda}{N}\Big),\quad\bar{u}_{i-1}:=\frac{\mathbb{E}S_{i-1}}{N}$
@@
$E\left[e^{\delta|X_{i}|}\,\big|\,u_{i-1}\in G_{i-1}\right]\leq K\quad\text{{(bounded MGF)}}$
@@
$\iff u_{i-1}\in G_{i-1}$
For the base case, we have $u_{0}\in G_{0}$ where
$G_{0}:=\Big(u_{0}-\frac{\lambda}{N},u_{0}+\frac{\lambda}{N}\Big);$
bounded MGF is satisfied for $i=1$. Thus concentration inequality holds,
@@
$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\geq\lambda\,\Big|\,u_{1}\in G_{1},...,u_{n-1}\in G_{n-1}\Big)\leq 2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
@@
$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\leq\lambda\,\Big|\,u_{1}\in G_{1},...,u_{n-1}\in G_{n-1}\Big)\geq 1-2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$

### v5

@@ Abstract
While classical concentration inequalities are typically restricted to two special cases—independence and martingale difference sequences—we extend concentration inequalities to a much broader class of stochastic processes by relaxing these foundational conditions. Specifically, in calculus terms, while independence and the martingale difference property correspond respectively to
@@
we relax these conditions to
@@
Furthermore, concentration inequalities for branching random walks are established.
@@ 1.1 Concentration inequalities
Let $(S_{i})_{i=1,2,...}$ be a real-valued random process, one-dimensional random walk. $S_{n}$ can be formulated as $S_{n}=X_{1}+X_{2}+...+X_{n}$; $X_{i}=S_{i}-S_{i-1}$. Under certain conditions, in particular independence, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT). For $S_{n}$ beyond the neighborhood, i.e., $|S_{n}-\mathbb{E}(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [24]), also referred as Chernoff bound,
$\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
@@
For a long time, there had been no substantive improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [42]. The reason might be two folds. First, there lacked motivation. Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results somehow concealed the essential reason why the inequalities hold.. In early 2000s [39] relaxed the boundedness to finite exponential moments; i.e. for a $\delta>0,\mathbb{E}(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda|)\leq e^{-c\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$— not Chernoff bound. Later, motivated by the study of the free energy of a directed polymer in a random environment, [34] gave Chernoff bound under the hypothesis that, for a $\delta>0,\mathbb{E}(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$). Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [26] [21]. However, to our knowledge, no existing work eliminates both the martingale and boundedness conditions while achieving sub-Gaussian concentration; related work by Wormald (1995) is discussed later. The first major contribution of this paper is to do so. Informally, the small per-step effect in a random process can be controlled by the Lipschitz continuity of the mean increment with respect to the history, not necessarily by “almost independence”.
1.2 Notation and contributions
@@
We let $\mathbb{E}_{i}[\,\cdot\,]$ abbreviate $\mathbb{E}[\,\cdot\mid u_{i}\,]$. Note that $\mathbb{E}_{i}[X_{t}]$ is a function of both $t$ and past state $(u_{i})_{i\leq t}$. To connect the calculus notation with the underlying stochastic process, we define
@@ 1.4 Motivation and contribution
In algorithm analysis and random graphs study, which are modelled as graph evolution stochastic processes (see e.g. [36], [44], [45], [46], [47] and references therein), the differential equation method (DEM) started in 80’s. Recently DEM has obtained some important results and received increasingly attention. The procedure of DEM (see [44] for a general framework, and [28] for its stronger version) consists of (a) verifying that the random parameters concerned concentrate around their means, and (b) finding solutions to the differential equations satisfied the means. Concentration inequalities (e.g. Heoffding-Azuma inequalities) are used to prove concentration phenomena in DEM.
@@
2 Extension of Hoeffding-Azuma inequality
As a pedagogical warm-up, we present a “downgraded” proof for Azuma’s inequality. To start with, we introduce a simple inequality following [34] and [43].
@@ 2 Extension of Hoeffding-Azuma inequality
$\Pr(|S_{n}-\mathbb{E}S_{n}|\geq\lambda)\leq 2\exp\Big(-\frac{\lambda^{2}}{2nc^{2}}\Big)$
@@
which, given $\lambda=o(n)$, is exponentially small, $\exp(-n^{p})$ with $0<p<1$, under the choice $t\sim n^{(p-1)/2}$.
Proof (A non-sharp version of Azuma’s inequality).
We only consider ${\Pr}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$, as the proof of the reverse ${\Pr}(S_{n}-\mathbb{E}S_{n}\leq-\lambda)$ is exactly the same. By Markov’s inequality, inductive application of Lemma 2.1 and optimization of $t$, we have
${\Pr}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$ $\leq e^{-t\lambda}\cdot E\Big[e^{tS_{n-1}}\cdot\mathbb{E}(e^{tX_{n}}\mid X_{1},...,X_{n-1})\Big]$
@@
$\leq e^{-t\lambda}\cdot\mathbb{E}\Big[e^{tS_{n-1}}\cdot e^{t^{2}e^{c}}\Big]~~~~~~~\mbox{(}|X_{i}|\leq c\mbox{)}$
@@
The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [34] and [43]) without compromising strength of results; from the practical point of view, the difference between the constant factors in the exponent, i.e., $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ versus $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$, is immaterial – both are square-exponential decay. In addition, it is shown that sub-Gaussian tail bounds can be obtained without “nifty” techniques.
Next, we show what weaker conditions can ensure concentration inequalities, beyond independence, martingale differences, and boundedness. To this end, define Doob’s (or McDiarmid’s) martingale
@@
$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\,,\quad E_{i-1}(d_{i})=0$
Note
@@
Note, $\mathbb{E}_{i}(X_{\ell})\equiv\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
@@
With this notation, we have
@@
$\leq\mathbb{E}_{X_{i}^{\prime}}\Big[\,\bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\bigl|\,\Big]$
@@
and $\mathbb{E}[\,e^{\delta A|X_{i}|}\mid u_{i-1}\,]\leq K$ (exponential moment existence) for a constant $\delta>0$ , $i=1,2,$..., then for $n\leq M$ $\dagger$$\dagger\daggerActually$, the lemma holds for $n=O(M)$; we assume $n\leq M$ here to simplify the argument.
$\Pr\big(\,|\,S_{n}-\mathbb{E}S_{n}\,|\geq\lambda\,\big)\leq 2\exp{\Big(\!-c\frac{\lambda^{2}}{n}\Big)},\quad c=\frac{\delta^{2}}{4K^{2}}\;$ (2.9)
Remarks.
    where
  • For $n=O(N)$, the L-Lipschitz condtion may imply
    $\left|\sum_{\ell=i+1}^{n}\mathbb{E}_{i-1}(X_{\ell}\mid X_{i})-\sum_{\ell=i+1}^{n}\mathbb{E}_{i-1}(X_{\ell}\mid X_{i}^{\prime})\right|\leq L\displaystyle{|X_{i}-X_{i}^{\prime}|}$
    When $X_{i}$ and $X_{i}^{\prime}$ are bounded, this reduces to McDiarmid’s bounded difference condition [30] when applied to random walks.
  • Independence and martingale difference assumptions both imply this L-Lipschitz condition (2.8). That is, the Lipschitz condition is weaker than those assumptions.
  • Poisson, Binomial, and Geometric distributions, for example, all satisfy the finite exponential moment condition, as does the Gaussian distribution.
@@
$\mathbb{E}_{i-1}\left[e^{\delta|d_{i}|}\right]=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
@@
${\Pr}(S_{n}-\mathbb{E}S_{n}\geq\lambda)\leq\exp\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)$
For the other side of the concentration inequality, we refer to the fact that Lemma 2.1 holds for negative $t$ as well. The proof is identical to the above. ∎
@@
  • The condition of Lipschitz continuity (2.10) may be interpreted as “continuity” of the trajectory with respect to the history; i.e. small differences in initial position do not lead to vehement differences in the trajectory.
@@
Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on. We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the domain in question lies sufficiently away from singularities).
@@
If $\mathbb{E}\big(e^{\delta A|X_{i}|}\mid u_{i-1}\big)\leq K$ (existence of an exponential moment), then in the domain where $\mu_{i}$ and $\nu_{i}$ are smooth, the concentration inequality holds.
@@
$\frac{\partial\mu(t,u)}{\partial u},\quad\frac{\partial^{2}\mu(t,u)}{\partial u^{2}},\quad\frac{\partial\nu(t,u)}{\partial u},\quad\cdots$
are uniformly bounded in the domain of smoothness (i.e., away from singularities), in addition to the exponential moment condition. These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE). From the PDE we have every partial derivatives of $\bar{u}$, and $u$ is close to $\bar{u}$ with exponentially rare and negligible exceptions. Note, the condition (2.10) says that small changes in distant past generations won’t lead to violently change in the current step size. Theorem 2.2 tells us, in terms of PDE of $\bar{u}$, what local behavior of $u$ can meet (2.10), for the concentration inequality to hold.
Proof of Theorem 2.2.
We only need to show Lipschitz continuity (2.10) holds, given the smoothness of $\mu_{i}$ and $\nu_{i}$. [^3] In most cases of interest, when $\mathbb{E}(X|u)$ is smooth, $\mathbb{E}(X^{2}|u)$ is smooth as well. Intuitively, the requirement that $\mathbb{E}(X^{2}|u)$ be smooth is the price we pay for not assuming boundedness of $X$. The proof is proceeded by induction. The base case holds, i.e. $|\mathbb{E}(X_{i+1}|u_{i})-\mathbb{E}(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
@@
the following recurrence equation about $\mathrm{Var}_{n+1}$ holds,
@@
Now let’s estimate $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$. From (2.13) and (2.12), we have
@@
Given above, it is easy to check that, in the domain where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$. Further, by induction, $\frac{d\mathrm{Var}_{n+1}}{du}$ exists and its order is $O(1/N)$.
@@ 3 Concentration inequalities in the context of BRW
$\mathbb{E}^{\!{}^{(M)}}[\,\cdot\,]=\frac{\displaystyle\int(\,\cdot\,)\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}{\displaystyle\int\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}$ (3.2)
@@
$Pr[\ |S_{n}-\mathbb{E}(S_{n})|\geq\lambda]\leq 2\exp{\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)},\quad\text{$n\leq M$ }$ (3.8)
where $\mathbb{E}(S_{n})$ is given by (3.2).
@@
If $E(e^{\delta A|X_{i}|}\mid u_{i-1})\leq K$, then in the domain where $m_{i}(u)$, $\mu_{i}(u)$ and $\nu_{i}(u)$ are smooth, the concentration inequality holds.
@@
$\overline{X}_{n+1}^{(k)}\equiv\mathbb{E}_{n}^{(k)}(X_{n+1}),\quad f_{n+1}^{(k-1)}(u_{n+1})\equiv\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right];$
@@
which is smooth in $u_{n}$. Observe that the term $\frac{\overline{X}_{n+1}^{(k)}}{N}{f_{n+1}^{(k-1)}}^{\prime}(u_{n})$ involves no second order derivatives of $f_{n+2}^{(k-2)}$ which is absorbed into the $O(N^{-2})$ term. Thus, we have proved that for any $k$, $\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]$ is smooth in $u_{n}$, where
@@
In the situation of bounded $X_{i}$, a particle always stays around the mean trajectory as long as its birthplace $u_{i-1}$ is near that trajectory. In the case of unboundedness, however, a particle may move into regions near singularities (e.g., critical points in K–SAT) where $u_{i}$ is not presumed smooth. In such regions, $\mathbb{E}(e^{\delta Xi})<K$ may fail, among others. The good news is, because concentration inequality holds in the neighborhood area of the mean trajectory—is at least $\epsilon$ far away from the singularity—the outliers account for a negligible proportion of the population and can be ignored. Crucially, pruning these rare extremal paths does not weaken the concentration. The following theorem formalizes this observation.
@@
Let $\epsilon$ be small compared with $\delta$, $N^{1/2}\lesssim\lambda<\epsilon N$, and as before,
@@
$D_{i-1}:=\Big(\bar{u}_{i-1}-\frac{\lambda}{N},\bar{u}_{i-1}+\frac{\lambda}{N}\Big),\quad\bar{u}_{i-1}:=\frac{\mathbb{E}S_{i-1}}{N}$
@@
$\mathbb{E}\left[e^{\delta A|X_{i}|}\,\big|\,u_{i-1}\in D_{i-1}\right]\leq K\quad\textnormal{{(bounded MGF)}}$
@@
$\iff u_{i-1}\in D_{i-1}$
For the base case, we have $u_{0}\in D_{0}$ where
$D_{0}:=\Big(u_{0}-\frac{\lambda}{N},u_{0}+\frac{\lambda}{N}\Big);$
bounded MGF is satisfied for $i=1$. Thus the concentration inequality holds as
@@
$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\geq\lambda\,\Big|\,u_{1}\in D_{1},...,u_{n-1}\in D_{n-1}\Big)\leq 2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
@@
$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\leq\lambda\,\Big|\,u_{1}\in D_{1},...,u_{n-1}\in D_{n-1}\Big)\geq 1-2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$